OpenStudy (anonymous):

Mr. Daton purchased men's coats for $80 each and sold them for $158 each. Find the markup on the retail to the nearest whole percent.

OpenStudy (anonymous):

49.3%

OpenStudy (anonymous):

So 49%

OpenStudy (anonymous):

How did you come to that answer?

OpenStudy (anonymous):

markup = retail price - cost price = 158 - 80 = $78 percentage = 78/158 * 100 = 49.3%

OpenStudy (anonymous):

Understand?

OpenStudy (anonymous):

Let me work through it on paper right quick and make sure

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Yes, I see now. Thank you so much for your help!!

OpenStudy (anonymous):

No problem!!!